OpenStudy (anonymous):

Evaluate: (4/9)^1/2 A) 2/3 B) 4/3 C) 4/9 D) 2/9

OpenStudy (anonymous):

What is the square root of 4 and 9 ??

OpenStudy (anonymous):

\[x ^{1/2}=\sqrt{x}\] so \[\sqrt{4}/\sqrt{9}\]

OpenStudy (anonymous):

The square root of: 4=2 9=3

OpenStudy (anonymous):